# Through the Dual-Prism: A Spectral Perspective on Graph Data Augmentation for Graph Classification

### v1

@@ Abstract
Graph Neural Networks (GNNs) have become the preferred tool to process graph data, with their efficacy being boosted through graph data augmentation techniques. Despite the evolution of augmentation methods, issues like graph property distortions and restricted structural changes persist. This leads to the question: Is it possible to develop more property-conserving and structure-sensitive augmentation methods? Through a spectral lens, we investigate the interplay between graph properties, their augmentation, and their spectral behavior, and found that keeping the low-frequency eigenvalues unchanged can preserve the critical properties at a large scale when generating augmented graphs. These observations inform our introduction of the Dual-Prism (DP) augmentation method, comprising DP-Noise and DP-Mask, which adeptly retains essential graph properties while diversifying augmented graphs. Extensive experiments validate the efficiency of our approach, providing a new and promising direction for graph data augmentation.
@@
11footnotetext: Y. Xia and R. Yu contributed equally to this work.22footnotetext: Y. Liang and X. Wang are the corresponding authors.
1 Introduction
@@
Graph structures, modeling complex systems through nodes and edges, are ubiquitous across various domains, including social networks (Newman et al., 2002), bioinformatics (Yi et al., 2022), and transportation systems (Jin et al., 2023a). Graph Neural Networks (GNNs) (Kipf & Welling, 2016a) elegantly handle this relational information, paving the way for tasks such as accurate predictions. Their capabilities are further enhanced by graph data augmentation techniques. These methods artificially diversify the dataset through strategic manipulations, thereby bolstering the performance and generalization of GNNs (Rong et al., 2019; Feng et al., 2020; You et al., 2020). Graph data augmentation has progressed from early random topological modifications, exemplified by DropEdge (Rong et al., 2019) and DropNode (Feng et al., 2020), to sophisticated learning-centric approaches like InfoMin (Suresh et al., 2021). Furthermore, techniques inspired by image augmentation’s mixup principle (Zhang et al., 2017) have emerged as prominent contenders in this domain (Verma et al., 2019; Wang et al., 2021; Guo & Mao, 2021).
@@
Though promising, these augmentation methods are challenged by three key issues as follows. (1) Graph Property Distortion. Before the era of deep learning, graph properties, e.g., graph connectivity and diameter, served as vital features for classification for decades (Childs et al., 2009). While now they seem to be ignored, many aforementioned contemporary augmentation methods appear to sidestep this tradition and overlook the graph properties. For instance, an example graph from the IMDB-BINARY dataset (Morris et al., 2020) and its augmented graph via DropEdge are illustrated in Figures 1a and 1b, respectively. The polar plot in Figure 1e shows the properties of these graphs, where each axis represents a distinct property. It is evident that DropEdge significantly alters the original graph’s properties, as indicated by the stark difference between the shapes of the orange (original) and blue (augmented) pentagons. (2) Limited Structural Impact. The majority of existing methods’ localized alterations do not capture the broader relationships and structures within the graph, limiting their utility. Consider a social network graph, where removing an edge affects just the immediate node and does little to alter the overall community structure. We thus ask: Can we design more property-retentive and structure-aware data augmentation techniques for GNNs?
Through the Dual-Prism: A Spectral Lens. Graph data augmentation involves altering components of an original graph. These modifications, in turn, lead to changes in the graph’s spectral frequencies (Ortega et al., 2018). Recent research highlighted the importance of the graph spectrum: it can reveal critical graph properties, e.g., connectivity and radius (Chung, 1997; Lee et al., 2014). Additionally, it also provides a holistic summary of a graph’s intrinsic structure (Chang et al., 2021), providing a global view for graph topology alterations. Building on this foundation, a pivotal question arises: Could the spectral domain be the stage for structure-aware and property-retentive augmentation efforts? Drawing inspiration from dual prisms—which filter and reconstruct light based on spectral elements—can we design a polarizer to shed new light on this challenge? With this in mind, we use spectral graph theory, aiming to answer the following questions: 1) Can a spectral approach to graph data augmentation preserve essential graph properties effectively? 2) How does spectral-based augmentation impact broader graph structures? 3) How does spectral-based augmentation compare to existing methods in enhancing the efficiency of GNNs for graph classification?
We begin with an empirical exploration in Section 3, where we aim to understand the interplay between topological modifications and their spectral responses. Our insights reveal that changes in graph properties mainly manifest in low-frequency components. Armed with this, in Section 4, we unveil our Dual-Prism (DP) augmentation strategies, DP-Noise and DP-Mask, by only changing the high-frequency part of the spectrum of graphs. Figures 1c and 1d provide a visualization of the augmented graphs via our proposed methods, i.e., DP-Noise and DP-Mask. As shown in Figure 1e, compared with DropEdge, our approaches skillfully maintain the inherent properties of the original graph, differing only slightly in the ASPL. Note that although we solely present one example underscoring our method’s capability, its robustness is consistently evident across all scenarios.
In addition to the properties, we further explore the spectrum comparison, shown in Figure 1f. Compared with DropEdge, the spectrum shifts caused by our methods are noticeably smaller. Interestingly, despite our approaches’ relative stability in the spectral domain, they induce substantial changes in the spatial realm (i.e., notable edge modifications). This spectral stability helps retain the core properties, while the spatial variations ensure a rich diversity in augmented graphs. Conversely, DropEdge, despite only causing certain edge changes, disrupts the spectrum and essential graph properties significantly. Simply put, our methods skillfully maintain graph properties while also diversifying augmented graphs. In Section 5, we evaluate the efficacy of our methods on graph classification, across diverse settings: supervised, semi-supervised, unsupervised, and transfer learning on various real-world datasets. Our concluding thoughts are presented in Section 6.
Contributions. Our main contributions are outlined as follows. (1) Prism – Bridging Spatial and Spectral Domains: We introduce a spectral lens to shed light on spatial graph data augmentation, aiming to better understand the spectral behavior of graph modifications and their interplay with inherent graph properties. (2) Polarizer – Innovative Augmentation Method: We propose the globally-aware and property-retentive augmentation methods, Dual-Prism (DP), including DP-Noise and DP-Mask. Our methods can preserve inherent graph properties while simultaneously enhancing the diversity of augmented graphs. (3) New Light – Extensive Evaluations: We conduct comprehensive experiments spanning supervised, semi-supervised, unsupervised, and transfer learning paradigms on 21 real-world datasets. The experimental results demonstrate that our proposed methods can achieve state-of-art performance on the majority of datasets.
@@
Figure 1: Visualization of (a) a graph from the IMDB-BINARY dataset and its augmented graphs via (b) DropEdge (Rong et al., 2019), (c) DP-Noise (ours), and (d) DP-Mask (ours). Dashed line: Dropped edge. Red line: Added edge. (e) Five properties of these graphs. $r$: radius. $d$: diameter. conn.: connectivity. ASPL: average shortest path length. #peri: number of periphery. Ori.: Original. D.E.: DropEdge. DP-N: DP-Noise. DP-M: DP-Mask. (f) The eigenvalues of these four graphs.
2 Related work
Spectrum and GNNs. Spectral graph theory (Chung, 1997) has found an appealing intersection with GNNs (Ortega et al., 2018; Wu et al., 2019; Dong et al., 2020; Bo et al., 2021; Chang et al., 2021; Yang et al., 2022). Early GNN approaches employed the spectrum of the Laplacian matrix to define convolution operations on graphs in the spectral domain (Hammond et al., 2011; Defferrard et al., 2016). As it evolved, there was a strategic shift towards spectral methods to enhance scalability (Nt & Maehara, 2019). This spectral perspective continues to be influential across various graph learning domains, notably in graph contrastive learning (GCL) (Liu et al., 2022; Lin et al., 2022), adversarial attacks (Entezari et al., 2020; Chang et al., 2021), and multivariate time series learning (Cao et al., 2020; Jin et al., 2023b). Zooming into the GCL domain, where data augmentation plays a pivotal role, Liu et al. (2022) introduced the general rule of effective augmented graphs in GCL via a spectral perspective. Lin et al. (2022) presented a novel augmentation method for GCL, focusing on the invariance of graph representation in the spectral domain. Notably, while these studies offer valuable insights, they mainly concentrate on the GCL paradigm, often neglecting broader discussions about preserving the core properties of graphs and supervised tasks.
@@ 2 Related work
Data Augmentations for GNNs. Graph data augmentation refers to the process of modifying a graph to enhance or diversify the information contained within, which can be used to bolster the training dataset for better generalization or model variations in real-world networks Ding et al. (2022); Zhao et al. (2022). Early methods are grounded in random modification to the graph topology. Techniques like DropEdge (Rong et al., 2019), DropNode (Feng et al., 2020), and random subgraph sampling (You et al., 2020) introduce stochastic perturbations in the graph structure. In addition to random modification, there is a wave of methods utilizing more sophisticated, learning-based strategies to generate augmented graphs (Suresh et al., 2021). Another research line is inspired by the efficiency of mixup (Zhang et al., 2017) in image augmentation, blending node features or entire subgraphs to create hybrid graph structures (Verma et al., 2019; Wang et al., 2021; Guo & Mao, 2021; Han et al., 2022; Park et al., 2022; Ling et al., 2023). However, while the above techniques, either spatial or spectral, have advanced the field of graph data augmentation, challenges remain, especially in preserving broader structural changes and graph semantics. Our work presents a new spectral augmentation approach, providing a principled way to modulate a graph’s spectrum while preserving the core, low-frequency patterns.
3 A Spectral Lens on Graph Data Augmentations
3.1 Preliminaries
An undirected graph $\mathcal{G}$ is represented as $\mathcal{G}=(V,E)$ where $V$ is the set of nodes with $|V|=N$ and $E\subseteq V\times V$ is the set of edges. Let $A\in\mathbb{R}^{N\times N}$ be the adjacency matrix of $\mathcal{G}$, with elements $a_{ij}=1$ if there is an edge between nodes $i$ and $j$, and $a_{ij}=0$ otherwise. Let $D\in\mathbb{R}^{N\times N}$ be the degree matrix, which is a diagonal matrix with elements $d_{ii}=\sum_{j}a_{ij}$, representing the degree of node $i$. The Laplacian matrix of $\mathcal{G}$ is denoted as $L=D-A\in\mathbb{R}^{N\times N}$. The eigen-decomposition of $L$ is denoted as $U\Lambda U^{\top}$, where $\Lambda=diag(\lambda_{1},\dots,\lambda_{N})$ and $U=[u_{1}^{\top},\dots,u_{N}^{\top}]\in\mathbb{R}^{N\times N}$. For graph $\mathcal{G}$, $L$ has $n$ non-negative real eigenvalues $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{N}$. Specifically, the low-frequency components refer to the eigenvalues closer to 0, and the high-frequency components refer to the relatively larger eigenvalues.
3.2 Spectral Analysis Insights
Adopting a spectral viewpoint, we conduct a thorough empirical study to understand the interplay between graph properties, graph topology alterations in the spatial domain, and their corresponding impacts in the spectral realm. The findings from our analysis include three crucial aspects as detailed below. Details of experimental settings and more results can be found in Appendices C & D.
@@ 3.2 Spectral Analysis Insights
Obs 1. The position of the edge flip influences the magnitude of spectral changes.
@@
In Figures 2a and 2b, we explore how adding different edges to a toy graph affects its eigenvalues. For instance, the addition of the edge 1$\leftrightarrow$3 (shown as the red line), which connects two proximate nodes, primarily impacts the high-frequency component $\lambda_{6}$ (highlighted by the red rectangle). In contrast, when adding edge 2$\leftrightarrow$6 (the blue line) between two distant nodes, the low-frequency component $\lambda_{1}$ exhibits the most noticeable change (indicated by the blue rectangle). These variations in the spectrum underscore the significance of the edge-flipping position within the graph’s overall topology, whose insight is consistent with findings by (Entezari et al., 2020; Chang et al., 2021). Such spectral changes not only affect the graph’s inherent structural features but also potentially affect the outcomes of tasks relying on spectral properties, such as graph-based learning.
Obs 2. Low-frequency components display greater resilience to edge alterations.
@@
Building on Obs 1, we further investigate the phenomenon of different responses of high- and low-frequency components to topology alterations using a real-world graph from the REDDIT-BINARY dataset. We apply DropEdge (Rong et al., 2019) for data augmentation by first randomly dropping $20\%$ and $50\%$ of edges and then computing the corresponding eigenvalues, as depicted in Figure 2c. Our findings indicate that, under random edge removal, low-frequency components exhibit greater robustness compared to their high-frequency counterparts.
@@
Obs 3. Graph properties are crucial for graph classification.
Certain fundamental properties of graphs, e.g., diameter and radius, are critical for a variety of downstream tasks, including graph classification (Feragen et al., 2013). In Figures 3a and 3b, we present the distributions of two key graph properties – diameter $d$ and radius $r$ – across the two classes in the REDD-M12 dataset. The different variations in these distributions emphasize their critical role in graph classification. Nevertheless, arbitrary modifications to the graph’s topology, e.g., random-manner-based augmentation techniques, could potentially distort these vital properties, illustrated in Figures 1b and 1e.
Obs 4. Specific low-frequency eigenvalues are closely tied to crucial graph properties. From Obs 3, a question is raised: Can we retain the integrity of these essential graph properties during augmentation? To explore this, we turn our attention back to the toy graph in Figure 2a and examine the evolution of its properties and eigenvalues in response to single-edge flips. In Figures 3c and 3d, we chart the graph’s average shortest path length (denoted by blue dots), diameter $d$ (denoted by green dots) against the reciprocal of its second smallest eigenvalue $1/\lambda_{1}$ (denoted by red dots)***For visual clarity, we scaled $d$ and $1/\lambda_{1}$ by dividing it by 3 and 5, respectively.. Our observations reveal a notable correlation between the alterations in $d$ and $1/\lambda_{1}$.
Further, we investigate correlations among overall spectral shifts, graph properties, and specific eigenvalues. Consistent with established methodologies (Lin et al., 2022; Wang et al., 2022; Wills & Meyer, 2020), we adopt the Frobenius distance to quantify the overall spectral variations by computing the $L_{2}$ distance between the spectrum of $\mathcal{G}$ and augmented graph $\mathcal{G}^{\prime}$. Notably, this spectral shift does not directly correspond with the changes in properties or eigenvalues. This suggests that by maintaining critical eigenvalues, primarily the low-frequency components, we can inject relatively large spectral changes without affecting essential graph properties. This observation thus leads to our proposition: preserving key eigenvalues while modifying others enables the generation of augmented graphs that uphold foundational properties, instead of only focusing on the overall spectral shifts.
4 Methodology
Drawing inspiration from how prisms decompose and reconstruct light and how a polarizer selectively filters light (see Figure 4a), we design our own “polarizer”, i.e., the Dual-Prism (DP) method for graph data augmentation, as depicted in Figure 4b. The details are illustrated below, followed by both empirical evidence and theoretical rationale to substantiate our approach.
@@ 4 Methodology
4.1 Proposed Augmentation Methods
The proposed Dual-Prism (DP) methods, including DP-Noise and DP-Mask, obtain the augmented graphs by directly changing the spectrum of graphs. A step-by-step breakdown is delineated in Algorithm 1. The DP method starts by extracting the Laplacian Matrix $L$ of a graph $\mathcal{G}$ by $L=D-A$ and then computes its eigen-decomposition $L=U\Lambda U^{\top}$. Based on the frequency ratio $r_{f}$, $N_{a}$ eigenvalues are selected for augmentation, where $N_{a}=N\times r_{f}$. Note that since we only target high-frequency eigenvalues, we arrange eigenvalues in increasing order and only focus on the last $N_{a}$ eigenvalues. Then, a binary mask $M$ is formed based on the augmentation ratio $r_{a}$ to sample the eigenvalues to make a change. Depending on the chosen augmentation type $T$, either noise is infused to the sampled eigenvalues, modulated by $\sigma$ and $M$ (from Line 6 to 8), or the eigenvalues are adjusted using the mask $M$ directly (from Line 9 to 10). Finally, we reconstruct the new Laplacian $\hat{L}$ based on the updated eigenvalues $\hat{\Lambda}$. Given the Laplacian matrix is $L=D-A$, where $D$ is a diagonal matrix, an updated adjacency matrix $\hat{A}$ can be derived by eliminating self-loops. Lastly, we obtain the augmented graph $\hat{\mathcal{G}}$ with its original labels and features retained.
@@ 4.1 Proposed Augmentation Methods
4:$U$ and $\Lambda\leftarrow$ eigenvalue decomposition of $L$triangleright $\Lambda$ is arranged in increasing order.
5:$N_{a}\leftarrow int(N\times r_{f})$triangleright Get the number of eigenvalues to be augmented.
@@
9:     for $i\in\{1,\cdots,N_{a}\}$ do $\{\lambda_{N\!-\!i}\leftarrow\max(0,\lambda_{N\!-\!i}+\sigma M_{i}\epsilon_{i})\}$triangleright Add noise to the high-frequency part.
@@
11:     for $i\in\{1,\cdots,N_{a}\}$ do $\{\lambda_{N\!-\!i}\leftarrow(1-M_{i})\lambda_{i}\}$$\triangleright$ Mask the high-frequency part.
12:$\hat{L}\leftarrow U^{\top}\hat{\Lambda}U,\hat{A}\leftarrow\!-\!\hat{L}$triangleright Calculate the new Laplacian and new adjacent matrix.
@@
4.2 Empirical Evidence
Next, we aim to verify the correctness of our methods on the improvement of the performance for graph classification via experimental analysis.

### v3

@@ Abstract
Graph Neural Networks have become the preferred tool to process graph data, with their efficacy being boosted through graph data augmentation techniques. Despite the evolution of augmentation methods, issues like graph property distortions and restricted structural changes persist. This leads to the question: Is it possible to develop more property-conserving and structure-sensitive augmentation methods? Through a spectral lens, we investigate the interplay between graph properties, their augmentation, and their spectral behavior, and observe that keeping the low-frequency eigenvalues unchanged can preserve the critical properties at a large scale when generating augmented graphs. These observations inform our introduction of the Dual-Prism (DP) augmentation methods, including DP-Noise and DP-Mask, which retain essential graph properties while diversifying augmented graphs. Extensive experiments validate the efficiency of our approach, providing a new and promising direction for graph data augmentation.
Codehttps://github.com/yu-rp/DualPrism
@@
footnotetext: * Y. Xia and R. Yu contributed equally to this work.footnotetext: Y. Liang and X. Wang are the corresponding authors.
Introduction
@@
Graph structures, modeling complex systems through nodes and edges, are ubiquitous across various domains, including social networks (Newman, Watts, and Strogatz 2002), bioinformatics (Yi et al. 2022), and transportation systems (Jin et al. 2023a). Graph Neural Networks (GNNs) (Kipf and Welling 2016a) elegantly handle this relational information for tasks like accurate predictions. Their capabilities are further enhanced by graph data augmentation techniques, which artificially diversify the dataset via strategic manipulations, thereby improving the performance and generalization of GNNs. Graph data augmentation has progressed from early random topological modifications, e.g. DropEdge (Rong et al. 2019) and DropNode (Feng et al. 2020), to sophisticated learning-centric approaches like InfoMin (Suresh et al. 2021). Furthermore, techniques inspired by image augmentation’s mixup principle (Zhang et al. 2017) have emerged as prominent contenders (Verma et al. 2019; Wang et al. 2021; Guo and Mao 2021).
@@
Though promising, these augmentation methods are challenged by three key issues as follows. (1) Graph Property Distortion. Before the era of deep learning, graph properties, e.g., graph connectivity and diameter, served as vital features for classification for decades (Childs et al. 2009). While now they seem to be ignored, many aforementioned contemporary augmentation methods appear to sidestep this tradition and overlook the graph properties. For instance, an example graph from the IMDB-BINARY dataset (Morris et al. 2020) and its augmented graph via DropEdge are illustrated in Figures 1a and 1b, respectively. The polar plot in Figure 1e shows the properties of these graphs, where each axis represents a distinct property. It is evident that DropEdge significantly alters the original graph’s properties, as indicated by the stark difference between the shapes of the orange (original) and blue (augmented) pentagons. (2) Limited Structural Impact. The majority of existing methods’ localized alterations do not capture the broader relationships and structures within the graph, limiting their utility. Consider a social network graph, where removing an edge affects just the immediate node and does little to alter the overall community structure. We thus ask: Can we design more property-retentive and structure-aware data augmentation techniques for GNNs?
@@
Figure 1: Visualization of (a) a graph from the IMDB-BINARY dataset and its augmented graphs via (b) DropEdge (Rong et al. 2019), (c) DP-Noise (ours), and (d) DP-Mask (ours). Dashed line: Dropped edge. Red line: Added edge. (e) Five properties of these graphs. $r$: radius. $d$: diameter. conn.: connectivity. ASPL: average shortest path length. #peri: number of periphery. Ori.: Original. D.E.: DropEdge. DP-N: DP-Noise. DP-M: DP-Mask. (f) The eigenvalues of these four graphs.
Through the Dual-Prism: A Spectral Lens. Graph data augmentation involves altering components of an original graph. These modifications, in turn, lead to changes in the graph’s spectral frequencies (Ortega et al. 2018). Recent research highlighted the importance of the graph spectrum: it can reveal critical graph properties, e.g., connectivity and radius (Chung 1997; Lee, Gharan, and Trevisan 2014). Additionally, it also provides a holistic summary of a graph’s intrinsic structure (Chang et al. 2021), providing a global view for graph topology alterations. Building on this foundation, a pivotal question arises: Could the spectral domain be the stage for structure-aware and property-retentive augmentation efforts? Drawing inspiration from dual prisms—which filter and reconstruct light based on spectral elements—can we design a polarizer to shed new light on this challenge? With this in mind, we use spectral graph theory, aiming to answer the following questions: 1) Can a spectral approach to graph data augmentation preserve essential graph properties effectively? 2) How does spectral-based augmentation impact broader graph structures? 3) How does spectral-based augmentation compare to existing methods in enhancing the efficiency of GNNs for graph classification?
We begin with an empirical exploration, where we aim to understand the interplay between topological modifications and their spectral responses. Our insights reveal that changes in graph properties mainly manifest in low-frequency components. Armed with this, we unveil our Dual-Prism (DP) augmentation strategies, DP-Noise and DP-Mask, by only changing the high-frequency part of the spectrum of graphs. Figures 1c and 1d provide a visualization of the augmented graphs via our proposed methods, i.e., DP-Noise and DP-Mask. As shown in Figure 1e, compared with DropEdge, our approaches skillfully maintain the inherent properties of the original graph, differing only slightly in the ASPL. Note that although we solely present one example underscoring our method’s capability, its robustness is consistently evident across all scenarios.
In addition to the properties, we further explore the spectrum comparison, shown in Figure 1f. Compared with DropEdge, the spectrum shifts caused by our methods are noticeably smaller. Interestingly, despite our approaches’ relative stability in the spectral domain, they induce substantial changes in the spatial realm (i.e., notable edge modifications). This spectral stability helps retain the core properties, while the spatial variations ensure a rich diversity in augmented graphs. Conversely, DropEdge, despite only causing certain edge changes, disrupts the spectrum and essential graph properties significantly. Simply put, our methods skillfully maintain graph properties while also diversifying augmented graphs. In the Experiments section, we evaluate the efficacy of our methods on graph classification, across diverse settings: supervised, semi-supervised, unsupervised, and transfer learning on various real-world datasets. Our concluding thoughts are presented in the last section.
Contributions. Our main contributions are outlined as follows. (1) Prism – Bridging Spatial and Spectral Domains: We introduce a spectral lens to shed light on spatial graph data augmentation, aiming to better understand the spectral behavior of graph modifications and their interplay with inherent graph properties. (2) Polarizer – Innovative Augmentation Method: We propose the globally-aware and property-retentive augmentation methods, Dual-Prism (DP), including DP-Noise and DP-Mask. Our methods are able to preserve inherent graph properties while simultaneously enhancing the diversity of augmented graphs. (3) New Light – Extensive Evaluations: We conduct comprehensive experiments spanning supervised, semi-supervised, unsupervised, and transfer learning paradigms on 21 real-world datasets. The experimental results demonstrate that our proposed methods can achieve state-of-art performance on the majority of datasets.
Related Work
@@
Data Augmentations for GNNs. Graph data augmentation refers to the process of modifying a graph to enhance or diversify the information contained within, which can be used to bolster the training dataset for better generalization or model variations in real-world networks (Ding et al. 2022; Zhao et al. 2022). Early methods are grounded in random modification to the graph topology. Techniques like DropEdge (Rong et al. 2019), DropNode (Feng et al. 2020), and random subgraph sampling (You et al. 2020) introduce stochastic perturbations in the graph structure. In addition to random modification, there is a wave of methods utilizing more sophisticated, learning-based strategies to generate augmented graphs (Suresh et al. 2021). Another research line is inspired by the efficiency of mixup (Zhang et al. 2017) in image augmentation, blending node features or entire subgraphs to create hybrid graph structures (Verma et al. 2019; Wang et al. 2021; Guo and Mao 2021; Han et al. 2022; Park, Shim, and Yang 2022; Ling et al. 2023). However, while the above techniques have advanced the field of graph data augmentation, challenges remain, especially in preserving broader structural changes and graph semantics.
@@
Spectrum and GNNs. Spectral graph theory (Chung 1997) has significantly influenced GNNs(Ortega et al. 2018; Wu et al. 2019; Dong et al. 2020; Bo et al. 2021; Chang et al. 2021; Yang et al. 2022), evolving from defining early GNN convolutions using the Laplacian spectrum (Hammond, Vandergheynst, and Gribonval 2011; Defferrard, Bresson, and Vandergheynst 2016) to enhancing scalability (Nt and Maehara 2019). This approach is pivotal in various domains, including graph contrastive learning (GCL) (Liu et al. 2022; Lin, Chen, and Wang 2022; Yang et al. 2023; Chen, Lei, and Wei 2024), adversarial attacks (Entezari et al. 2020; Chang et al. 2021), and multivariate time series (Cao et al. 2020; Jin et al. 2023b). Zooming into the GCL domain, where data augmentation plays a pivotal role, (Liu et al. 2022) introduced the general rule of effective augmented graphs in GCL via a spectral perspective; GCL-SPAN (Lin, Chen, and Wang 2022) presented a novel augmentation method for GCL, focusing on the invariance of graph representation in the spectral domain; GASSER (Yang et al. 2023) uses a selective perturbation across various frequency bands to maintain homophily ratios while introducing perturbations to optimize task-relevant features; and PolyGCL (Chen, Lei, and Wei 2024) demonstrates the spectral filters can be effectively utilized for contrastive learning on graphs. Though promising, these methods do not fully address the intricacies of maintaining inherent graph properties during augmentation.
A Spectral Lens on Graph Data Augmentations
Preliminaries. An undirected graph $\mathcal{G}$ is represented as $\mathcal{G}=(V,E)$ where $V$ is the set of nodes with $|V|=N$ and $E\subseteq V\times V$ is the set of edges. Let $A\in\mathbb{R}^{N\times N}$ be the adjacency matrix of $\mathcal{G}$, with elements $a_{ij}=1$ if there is an edge between nodes $i$ and $j$, and $a_{ij}=0$ otherwise. Let $D\in\mathbb{R}^{N\times N}$ be the degree matrix, which is a diagonal matrix with elements $d_{ii}=\sum_{j}a_{ij}$, representing the degree of node $i$. The Laplacian matrix of $\mathcal{G}$ is denoted as $L=D-A\in\mathbb{R}^{N\times N}$. The eigen-decomposition of $L$ is denoted as $U\Lambda U^{\top}$, where $\Lambda=diag(\lambda_{1},\dots,\lambda_{N})$ and $U=[u_{1}^{\top},\dots,u_{N}^{\top}]\in\mathbb{R}^{N\times N}$. For graph $\mathcal{G}$, $L$ has $n$ non-negative real eigenvalues $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{N}$. Specifically, the low-frequency components refer to the eigenvalues closer to 0, and the high-frequency components refer to the relatively larger eigenvalues.
Spectral Analysis Insights
Adopting a spectral viewpoint, we conduct a thorough empirical study to understand the interplay between graph properties, graph topology alterations in the spatial domain, and their corresponding impacts in the spectral domain. The findings from our analysis include four crucial aspects as detailed below.
@@ A Spectral Lens on Graph Data Augmentations
Obs 1. The position of the edge flip influences the magnitude of spectral changes. In Figures 2a and 2b, we explore how adding different edges to a toy graph affects its eigenvalues. For instance, the addition of the edge 1$\leftrightarrow$3 (shown as the red line), which connects two proximate nodes, primarily impacts the high-frequency component $\lambda_{6}$ (the red rectangle). In contrast, when adding edge 2$\leftrightarrow$6 (the blue line) between two distant nodes, the low-frequency component $\lambda_{1}$ exhibits the most noticeable change (the blue rectangle). These variations in the spectrum underscore the significance of the edge-flipping position within the graph’s overall topology, whose insight is consistent with findings by (Entezari et al. 2020; Chang et al. 2021). Such spectral changes not only affect the graph’s inherent structural features but also potentially affect the outcomes of tasks relying on spectral properties, such as graph-based learning.
@@
Obs 2. Low-frequency components display greater resilience to edge alterations. Building on Obs 1, we further investigate the different responses of high- and low-frequency components to topology alterations using a real-world graph. We apply DropEdge (Rong et al. 2019) for augmentation by first randomly dropping $20\%$ and $50\%$ of edges and then computing the corresponding eigenvalues, as depicted in Figure 2c. It can be observed that, under random edge removal, low-frequency components exhibit greater robustness compared to their high-frequency counterparts.
Obs 3. Graph properties are crucial for graph classification. Certain fundamental properties of graphs, e.g., diameter and radius, are critical for a variety of downstream tasks, including graph classification (Feragen et al. 2013). In Figures 4a and 4b, we present the distributions of two key graph properties – diameter $d$ and radius $r$ – across the two classes in the REDD-M12 dataset. The different variations in these distributions emphasize their critical role in graph classification. Nevertheless, arbitrary modifications to the graph’s topology, e.g., random-manner-based augmentation techniques, could potentially distort these vital properties, illustrated in Figures 1b and 1e.
Obs 4. Specific low-frequency eigenvalues are closely tied to crucial graph properties. From Obs 3, a question is raised: Can we retain the integrity of these essential graph properties during augmentation? To explore this, we turn our attention back to the toy graph in Figure 2a and examine the evolution of its properties and eigenvalues in response to single-edge flips. In Figures 4c and 4d, we chart the graph’s average shortest path length (denoted by blue dots), diameter $d$ (denoted by green dots) against the reciprocal of its second smallest eigenvalue $1/\lambda_{1}$ (denoted by red dots)***For visual clarity, we scaled $d$ and $1/\lambda_{1}$ by dividing it by 3 and 5, respectively.. Our observations reveal a notable correlation between the alterations in $d$ and $1/\lambda_{1}$.
Figure 3: (a) Training loss of GIN model on REDDIT-BINARY when graphs are augmented by masking different eigenvalues. (b) Graph classification performance on IMDB-BINARY under various hyperparameters. The lines represent the average accuracy, while the shaded areas indicate the error margins.
Further, we investigate correlations among overall spectral shifts, graph properties, and specific eigenvalues. Consistent with established methodologies (Lin, Chen, and Wang 2022; Wang et al. 2022; Wills and Meyer 2020), we adopt the Frobenius distance to quantify the overall spectral variations by computing the $L_{2}$ distance between the spectrum of $\mathcal{G}$ and augmented graph $\mathcal{G}^{\prime}$. Notably, this spectral shift does not directly correspond with the changes in properties or eigenvalues. This suggests that by maintaining critical eigenvalues, primarily the low-frequency components, we can inject relatively large spectral changes without affecting essential graph properties. This observation thus leads to our proposition: preserving key eigenvalues while modifying others enables the generation of augmented graphs that uphold foundational properties, instead of only focusing on the overall spectral shifts.
@@
Methodology
Drawing inspiration from how prisms decompose and reconstruct light and how a polarizer selectively filters light (Figure 5a), we design our own “polarizer”, i.e., the Dual-Prism (DP) method for graph data augmentation (Figure 5b).
@@
Proposed Augmentation Methods
The proposed DP methods, including DP-Noise and DP-Mask, obtain the augmented graphs by directly changing the spectrum of graphs. A step-by-step breakdown is delineated in Algorithm 1. The DP method starts by extracting the Laplacian Matrix $L$ of a graph $\mathcal{G}$ by $L=D-A$ and then computes its eigen-decomposition $L=U\Lambda U^{\top}$. Based on the frequency ratio $r_{f}$, $N_{a}$ eigenvalues are selected for augmentation, where $N_{a}=N\times r_{f}$. Note that since we only target high-frequency eigenvalues, we arrange eigenvalues in increasing order and only focus on the last $N_{a}$ eigenvalues. Then, a binary mask $M$ is formed based on the augmentation ratio $r_{a}$ to sample the eigenvalues to make a change. Depending on the chosen augmentation type $T$, either noise is infused to the sampled eigenvalues, modulated by $\sigma$ and $M$ (from Line 6 to 8), or the eigenvalues are adjusted using the mask $M$ directly (from Line 9 to 10). Finally, we reconstruct the new Laplacian $\hat{L}$ based on the updated eigenvalues $\hat{\Lambda}$. Given the Laplacian matrix is $L=D-A$, where $D$ is a diagonal matrix, an updated adjacency matrix $\hat{A}$ can be derived by eliminating self-loops. Lastly, we obtain the augmented graph $\hat{\mathcal{G}}$ with its original labels and features retained.
@@
Empirical Evidence
Next, we verify our methods on the improvement of the performance for graph classification via experimental analysis.
@@
# $\Lambda$ is arranged in increasing order.
4:$U$ and $\Lambda\leftarrow$ eigenvalue decomposition of $L$
# Get the number of eigenvalues to be augmented.
5:$N_{a}\leftarrow int(N\times r_{f})$
@@
9:     for $i\in\{1,\cdots,N_{a}\}$ do$\{\lambda_{N\!-\!i}\!\!\leftarrow\!\!\max(0,\lambda_{N\!-\!i}+\sigma M_{i}\epsilon_{i})\}$
# Mask the high-frequency part.
@@
11:     for $i\in\{1,\cdots,N_{a}\}$ do $\{\lambda_{N\!-\!i}\leftarrow(1-M_{i})\lambda_{i}\}$
# Calculate the new Laplacian and new adjacent matrix.
12:$\hat{L}\leftarrow U^{\top}\hat{\Lambda}U,\hat{A}\leftarrow\!-\!\hat{L}$